OpenStudy (anonymous):

simplify 3x^ 4 y^ 2/ 2 xy

OpenStudy (anonymous):

3/2x^3y

OpenStudy (anonymous):

\[\frac {3x^ 4 y^ 2}{2 xy}= \frac{3}{2}x^3y \]

OpenStudy (anonymous):

yeah, same.. taks for clarification @dpasingh